### v2

# BrainFLORA: Uncovering Brain Concept Representation via Multimodal Neural Embeddings

Abstract.
Understanding how the brain represents visual information is a fundamental challenge in neuroscience and artificial intelligence. While AI-driven decoding of neural data has provided insights into the human visual system, integrating multimodal neuroimaging signals—such as EEG, MEG, and fMRI—remains a critical hurdle due to their inherent spatiotemporal misalignment. Current approaches often analyze these modalities in isolation, limiting a holistic view of neural representation. In this study, we introduce BrainFLORA, a unified framework for integrating cross-modal neuroimaging data to construct a shared neural representation. Our approach leverages multimodal large language models (MLLMs) augmented with modality-specific adapters and task decoders, achieving state-of-the-art performance in joint-subject visual retrieval task and has the potential to extend multitasking. Combining neuroimaging analysis methods, we further reveal how visual concept representations align across neural modalities and with real-world object perception. We demonstrate a fundamental alignment between the representational structures of visual objects as understood by AI and those found in the human brain, showing a shared framework for comprehending the physical world. Beyond methodological advancements, BrainFLORA offers novel implications for cognitive neuroscience and brain-computer interfaces (BCIs). Our code is available at https://github.com/ncclab-sustech/BrainFLORA.
Neuroscience, Neural Coding, Concept-Selective, Visual Decoding, BCI, Multimodality, Generative Model
Co-first authors: Dongyang Li and Haoyang Qin.
Corresponding authors: Chen Wei and Quanying Liu.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754996ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligenceccs: Cognitive science
1. Introduction
The brain processes external stimuli by receiving and encoding information through the coordinated activity of large-scale neural networks and intricate mechanisms. Recent advances, particularly in generative models and pre-trained techniques, have driven substantial progress in visual decoding from various neural modalities, including EEG (song2023decoding, ; li2024visual, ; zhang2024cognitioncapturer, ), fMRI (chen2023seeing, ; takagi2023high, ; scotti2024reconstructing, ; scottimindeye2, ), and MEG  (benchetrit2024brain, ; li2024visual, ). By leveraging contrastive learning and pre-trained learning, these methods are reshaping our capacity to interpret complex patterns of visual activity in the brain.
One core challenge in achieving advanced performance lies in establishing cross-modal representations for diverse neural data. Such data exhibit significant complexity and variability, influenced by factors such as individual subjects, acquisition devices, visual content, and spatial-temporal dynamics. Recent works such as NeuroBind (yang2024neurobind, ), which deploys distinct encoders for each type of neural data. Additionally, one-encoder architectures like UMBRAE (xia2024umbrae, ) have aimed to learn a uniform representation across subjects in fMRI for brain decoding tasks. While these efforts demonstrate progress in improving performance for various downstream tasks, they remain constrained by the inherent limitations of effective modal alignment. This challenge highlights the critical need for more cohesive and innovative frameworks that not only integrate multiple modalities but also facilitate each unique modal characteristic to enhance understanding of brain representations related to different neural signals.
Inspired by multimodal large language models (MLLMs) in processing input information in a wider variety of modalities (lu2024unified, ), we propose a framework to integrate multimodal neural data, to improve encoding and decoding performance and promote the understanding of visual representation in brain signal recordings from different neural modalities. In this work, we develop BrainFLORA, a multimodal model that processes three modalities (i.e., EEG, MEG and fMRI) within a unified alignment framework. BrainFLORA consists of multimodal encoders, and a trainable mixture of experts (MoE) that performs universal projection across modalities. Unlike previous approaches, the encoder in BrainFLORA is trained on three neural modalities, including 16,540 paired EEG-image samples (grootswagers2022human, ), 19,848 paired MEG-image samples, and 8,640 paired fMRI-image samples (hebart2023things, ), ensuring shared representations across all modalities. To evaluate the BrainFLORA model, we apply it for downstream tasks including retrieval, captioning, and reconstruction respectively.
Our work has three main contributions:
  • We propose a novel neural encoder to learn disentangled representations from neural data embeddings, enabling feature complementarity across subjects and heterogeneous neural modalities.
  • Our framework, BrainFLORA, enables seamless alignment and integration with various downstream tasks, significantly enhancing its adaptability and achieving SOTA retrieval performance (Tab. 1) in joint-subjects visual decoding benchmarks.
  • BrainFLORA is the first work to concurrently unify multiple neural modalities (EEG, MEG and fMRI) and uncover the unique contribution of each modal to concept representation in the brain.
Figure 1. Overview of BrainFLORA and the comparison between concept spaces of EEG, MEG, and fMRI. The framework links conceptual Representation Similarity Matrices (RSMs) that capture visual relationships among objects to neural responses elicited by diverse object images from the THINGS database.
2. Related Work
2.1. Visual Decoding via Different Neural Modalities
For the past few years, significant progress has been made in visual decoding through neural modalities such as EEG, fMRI, and MEG. Early research predominantly focused on leveraging deep learning to classify (schirrmeister2017deep, ; spampinato2019deeplearninghumanmind, ; zheng2021attention, ) and retrieve (ye2022seeselfsupervisedcrossmodalretrieval, ) visual stimuli based on neural recordings. In recent years, visual decoding has achieved unprecedented quality (sun2023contrastattenddiffusedecode, ; zhouclip, ; benchetrit2024brain, ; ma2024alignedllmnewmultimodal, ; li2024visual, ), driven by advancements in generative models (ho2020denoisingdiffusionprobabilisticmodels, ; rombach2022high, ; liu2022flowstraightfastlearning, ) and LLMs (touvron2023llamaopenefficientfoundation, ; brown2020languagemodelsfewshotlearners, ). For instance, RealMind (li2024realmind, ) utilizes multimodal models that combine the ability of semantic and geometric information, leading to improved decoding performance. Additionally, MindEye2 (scottimindeye2, ) demonstrates high-quality cross-subject reconstruction and retrieval capabilities with limited sample sizes. Moreover, ATM (li2024visual, ) successfully achieves real-time level retrieval and generation by capitalizing on the high temporal resolution of EEG. These advancements underscore the practicality and effectiveness of utilizing pre-trained models to gain insights into visual information decoding and the functioning of the visual system.
2.2. Human Concept Embeddings
A growing body of research has explored how human conceptual representations can be modeled through similarity-based decision tasks. Seminal work by Hebart et al. (hebart2020revealing, ) and Fu et al. (fu2023dreamsim, ) demonstrated that human perceptual judgments reveal structured low-dimensional embeddings of object concepts. Building on this foundation, Wei et al. (wei2024cocog, ) proposed CoCoG, a novel visual generation framework that operationalizes these concept embeddings as control variables, enabling both behavioral prediction and controllable image synthesis while providing insights into cognitive mechanisms.
Recent advances have further quantified the alignment between machine and human concept spaces. Through large-scale behavioral experiments involving millions of ”odd-one-out” triples, Du et al. (du2025human, ) derived a 66-dimensional embedding space for 1,854 objects, combining psychophysical data with neuroimaging to establish significant correspondences between human and Artificial Neural Networks (ANNs) concept representations. This line of work is supported by complementary studies (han2024investigating, ; muttenthaler2023improving, ) that systematically evaluate representational alignment across modalities, confirming that similarity-based paradigms effectively capture shared organizational principles between biological and artificial cognitive systems.
Figure 2. Overall Framework of BrainFLORA. BrainFLORA comprises neural modality encoders, a universal projection, and a Mixture of Experts (MoE) projection module, with separate outputs for each modality. Left: The modality-specific encoders transform the input signal into semantic tokens. Middle: The MoE projection module projects and aligns diverse neural modalities into a unified semantic representation space. Right: Various task heads facilitate different downstream tasks such as retrieval, captioning, and reconstruction. All modules are jointly trained during the same stage, optimizing computational efficiency.
2.3. Cross-Modal Alignment in MLLMs
Aligning semantic representation across modalities is crucial for MLLMs, which aim to process and integrate information from diverse sources like text, images, audio, and video. Key challenges in multimodal alignment include the heterogeneity of data representations, dimensionality mismatches, semantic inconsistencies across modalities, temporal dependencies in dynamic modalities, and sparsity of paired multimodal datasets. To address these challenges, several modules and techniques have been developed for data alignment in multimodal models, such as Adapters (sung2022vl, ; hu2022lora, ; gao2021clipadapterbettervisionlanguagemodels, ), Perceiver Resampler (gong2023multimodalgptvisionlanguagemodel, ; du2023jointmusiclanguageattention, ; xue2024xgenmmblip3familyopen, ), and Q-Former (huang2019ccnet, ; Wei_2020_CVPR, ; zhu2023minigpt4enhancingvisionlanguageunderstanding, ). Specifically, adapters used in VL-Adapter  (sung2022vl, ), LoRA (hu2022lora, ), and CLIP-Adapter  (gao2021clipadapterbettervisionlanguagemodels, ) allow models to align multimodal data by selectively adapting their representations for specific tasks without requiring full retraining. PaLM2-VAdapter(xiao2024palm2vadapterprogressivelyalignedlanguage, ) introduced lightweight adapter modules, enabling to selectively adapt its representations for specific tasks without requiring full retraining of the massive PaLM2 model. Perceiver Resampler used in models like MultiModal-GPT (gong2023multimodalgptvisionlanguagemodel, ), Joint Music-Language Attention (du2023jointmusiclanguageattention, ), and X-GenMM (xue2024xgenmmblip3familyopen, ) maps high-dimensional, variable-length data into a compact, fixed-size representation suitable for integration. Similarly, Flamingo(alayrac2022flamingo, ) adapted high-dimensional, variable-length visual data into a consistent set of fixed-size tokens. In the domain of neural data processing, NeuroBind  (yang2024neurobind, ) integrates neural data using modality-specific encoders and representation learning with images. The ongoing development of alignment techniques underscores the growing importance of bridging the gap between different modalities, paving the way for more powerful multimodal models in neural data processing.
3. Method
Formulation. Let $T$ represent the length of neural data, $C$ the number of channels, and $N$ the total number of data samples. Our objective is to derive neural embeddings $Z=\mathcal{E}(X)\in\mathbb{R}^{N\times F}$ from the brain activity data $X\in\mathbb{R}^{N\times C\times T}$ and when $X$ is used as fMRI modality, the dimension $T$ is not included, where $\mathcal{E}$ is the unified projector and $F$ is the projection dimension of the embeddings. Concurrently, we use the CLIP model to extract image embeddings $\hat{Z}\in\mathbb{R}^{N\times F}$ from images $I$. Our goal is to effectively align multimodal neural representations with image representations, as illustrated in Fig. 2.
3.1. Model Architecture
Our method aims to process multiple modalities of neural data via one unified framework. All cross-modal neural features are efficiently aligned through the MoE module, maximizing complementary strengths across modalities. Multi-stage training strategies are applied to the encoding and alignment procedure. In the training phase, encoders are trained with neural data and image pairs using a contrastive learning framework. Then each modality is aligned through the Unified Projector using pre-trained modality-specific encoders. During inference, neural embeddings from the Unified Projector can be used for a variety of zero-shot tasks, including EEG/MEG/fMRI image classification, retrieval, and image reconstruction.
3.1.1. Neural Feature Extraction Module.
This module is mainly employed to address the cross-subject challenges in multimodal neural data decoding. The core difficulty arises from the heterogeneous neural responses of different subjects to identical visual stimuli, leading to substantial inter-subject variability in the data. By incorporating multi-granular approaches, as proposed in (wang2024medformer, ), into the encoder architecture, inter-subject heterogeneity can be mitigated, thereby facilitating the learning of shared joint embeddings across subjects.
For input neural signal $X\in\mathbb{R}^{T\times C}$, the architecture first performs Multi-Granularity Time Patching with exponentially increasing patch lengths $\{2^{1},2^{2},...,2^{n}\}$: $x_{p}^{(i)}\in\mathbb{R}^{N_{i}\times(2^{i}\cdot C)},N_{i}=\lceil T/2^{i}\rceil$. Then we get a group tokens $x^{(i)}\in\mathbb{R}^{N_{i}\times D}$ by:
(1) $x^{(i)}=x_{p}^{(i)}W^{(i)}+W_{pos}[1:N_{i}]+W_{gr}^{(i)}$
where $W_{\text{pos}}[1:N_{i}]\in\mathbb{R}^{N_{i}\times D}$is the first $N_{i}$ rows of the positional embedding $W_{\text{pos}}$, and a learnable granularity embedding $W_{\text{gr}}^{(i)}\in\mathbb{R}^{1\times D}$. And the router $u^{(i)}$ is used in the multi-granularity self-attention (as described later) for each granularity:
(2) $u^{(i)}=W_{pos}[N_{i}+1]+W_{gr}^{(i)}$
The embeddings undergo a two-stage Multi-Granularity Attention mechanism. In the first stage, intra-granularity attention captures temporal dependencies within each scale by concatenating patch embeddings with their corresponding router. The intermediate sequence of embeddings $z^{(i)}\in\mathbb{R}^{(N_{i}+1)\times D}$ is formed as:
(3) $z^{(i)}=\text{Concat}(x^{(i)},u^{(i)})$
where $x^{(i)}\in\mathbb{R}^{N_{i}\times D}$ represents patch embeddings at scale $i$, and $u^{(i)}\in\mathbb{R}^{1\times D}$ is the router embedding. Self-attention is applied on each $z^{(i)}$ to update both $x^{(i)}$ and $u^{(i)}$:
(4) $x^{(i)}\leftarrow\text{Attn}_{\text{tra}}\left(x^{(i)},z^{(i)},z^{(i)}\right)$
(5) $u^{(i)}\leftarrow\text{Attn}_{\text{tra}}\left(u^{(i)},z^{(i)},z^{(i)}\right)$
Then, inter-granularity attention facilitates information exchange across scales through router interactions. All router embeddings are aggregated:
(6) $U=\text{Concat}(u^{(1)},u^{(2)},...,u^{(n)})\in\mathbb{R}^{n\times D}$
These router embeddings serve as query vectors to attend to features across scales to get $x^{(i)}=\sum_{k=1}^{n}U_{k}\cdot x^{(i)}$. Then the attended features are concatenated as:
(7) $H_{\text{attn}}=\text{Concat}(x^{(1)},x^{(2)},...,x^{(n)})\in\mathbb{R}^{(\sum_{i=1}^{n}N_{i})\times D}$
Finally, $H_{\text{attn}}$ is processed through a temporal-spatial convolution module to obtain $H_{\text{conv}}$, followed by an MLP to get the final encoding $Y\in\mathbb{R}^{D}$, which aligns with CLIP embeddings of the corresponding visual stimuli image.
3.1.2. Universal Projection Module.
Aligning the multimodal neural representation space with the visual embedding space of the CLIP model is a crucial step, particularly given the complexity and heterogeneity of the multimodal neural imaging-image pairs. We employ a MoE module (zhu2024uni, ), which has $K$ projection experts $E_{1},E_{2},$ $\dots,E_{K}$, where each expert is a two-layer MLP, and a soft router $R_{\text{soft}}$ to control the contribution of each expert.
Using a small network $g$, we use the soft router computes a score vector for each token, with a length equal to the number of experts $K$. Finally, a Softmax function is applied to obtain the final routing vector, yielding the final routing vector. It can be formulated as:
(8) $R_{\text{soft}}(x_{i})=\frac{Sigmoid(g(x_{i}))}{\sum Sigmoid(g(x_{i}))}$
To alleviate the heterogeneity of different neural modalities. $R_{\text{soft}}$ is a lightweight MLP designed to receive the inputs of different encoders, and calculate the routing weights $W_{\text{soft}}\in\mathbb{R}^{N\times F\times K}$ of each expert for each neural token, which can be formulated as:
(9) $W_{\text{soft}}(Z)=\sigma\cdot R_{\text{soft}}\left(Z\right)$
where $\sigma$ is the SoftMax function. Then we can obtain aligned neural tokens $Z\in\mathbb{R}^{N\times F}$ through a weighted sum of all experts’ output as follows:
(10) $Z=\sum_{k=1}^{K}W_{\text{soft},k}\cdot E_{k}(Z)$
where $w_{\text{soft},k}$ denotes the routing weight of the $k$-th projection expert.
Figure 3. Architecture of the neural feature extraction module. The original neural sequences from multiple variates are simultaneously embedded into tokens. Multi-granularity attention is applied to the embedded tokens of correlated variables, enhancing electrode-level correlations. The representations for each token are then assigned through the router layers. Subsequently, Temporal-Spatial convolution is employed to mitigate overfitting while improving the model’s capacity for temporal-spatial representation learning.
3.2. Training Objective
By leveraging neural data $X$ and corresponding image data $I$, BrainFLORA could utilize the potential of each neural modalities and align them in semantic representation space. Training consists of three progressive stages: visual-neural data contrast training, cross-modal neural alignment via Universal Projection and task-specific fine-tuning. We employed multiple approaches to loss functions to adapt to various downstream tasks.
3.2.1. High-level Visual Modeling.
This pipeline is primarily used to learn high level unified representation from the universal projector. We use the contrastive learning functions inspired by  (scotti2024reconstructing, ) to train both the multi-tower encoder and Universal Projection:
(11) $\mathcal{L}_{\text{SoftCLIP}}=-\sum_{i,j=1}^{N}\frac{\exp\left(\frac{t_{i}\cdot t_{j}}{\tau}\right)}{Z_{i}}\cdot\log\left(\frac{\exp\left(\frac{p_{i}\cdot t_{j}}{\tau}\right)}{Z_{i}^{(p)}}\right)$
(12) $Z_{i}=\sum_{m=1}^{N}\exp\left(\frac{t_{i}\cdot t_{m}}{\tau}\right),\quad Z_{i}^{(p)}=\sum_{m=1}^{N}\exp\left(\frac{p_{i}\cdot t_{m}}{\tau}\right)$
Where $p_{i}=\mathcal{E}(x_{i}),t_{i}=\text{CLIP}_{\text{Image}}(y_{i})$, $\tau$ is a temperature hyperparameter, and $x_{i}$ and $y_{i}$ represent the ithi-th neural data sample and image respectively. For generative tasks, including captioning and reconstruction, we incorporate Mean Squared Error (MSE) loss to ensure consistent model training for regression-based tasks. As a result, the overall loss function of BrainFLORA is formulated as a composite of these individual loss components. Inspired by MindEye (scotti2024reconstructing, ), the optimization of model parameters $\theta$ is then governed by the following objective function:
[…]
where $\alpha$ and $\beta$ are hyperparameters to control the balance of each loss function, and $\mathcal{L}_{\text{prior}}$ is a ’diffusion prior’ similar to UnCLIP(ramesh2022hierarchical, ).
3.2.2. Low-level Visual Modeling.
Previous works (scottimindeye2, ; li2024visual, ; zhang2024cognitioncapturer, ) have extensively used prior diffusion to enhance the transformation of neural data embeddings to image embedding distributions. We continue this approach with prior diffusion, but use the neural embeddings directly output by the Universal Projection module for retrieval evaluation—rather than after image prior enhancement, because this can more objectively measure the degree of alignment of neural data with the neural manifold of the image without over-reliance on image priors.
To train the low-level pipeline, the original universal projector goes through additional CNN upsampler to get $\hat{f}$, the approximation of original image feature map $f\in\mathbb{R}^{h\times w\times c}$. Then a new image $\hat{I}$ is reconstructed using the Stable Diffusion XL (podell2023sdxl, ) VAE decoder $D(\cdot)$ given $\hat{f}$, and a compound loss $\mathcal{L}$ is minimized:
(14) $)\mathcal{L_{\text{lowlevel}}}=\|D(f)-D(\hat{f})\|+\|f-\hat{f}\|+\lambda_{P}\mathcal{L}_{P}(D(\hat{f}$))
where $\mathcal{L}_{P}(\cdot)$ is a perceptual loss such as LPIPS (zhang2018unreasonable, ), and $\lambda_{P}$ is loss weights.
4. Experiment
4.1. Dataset and Implementation
4.1.1. Dataset.
We jointly trained BrainFLORA with the THINGS-EEG2 (gifford2022large, ), THINGS-MEG, and THINGS-fMRI (hebart2023things, ) datasets to achieve a unified semantic representation. Notably, we noticed that relatively few decoding studies (xue2024hybrid, ) have utilized the THINGS-EEG1 dataset (grootswagers2022human, ), which might be due to its low signal-to-noise ratio. This limitation is primarily attributed to two methodological factors: (i) the short presentation time of visual stimulus (50ms), and (ii) the lack of trial repetition (each stimulus viewed only once by each participant). Finally, we explored fine-tuning experiments using THINGS EEG1 in the Appendix. The THINGS-EEG2 training set consists of 10 subjects, 1654 concepts, each associated with 10 images, with 4 repetitions per image, while the testing set includes 200 concepts, each represented by a single image and 80 repeated trials. The original training set for THINGS-MEG contains 4 subjects, 1854 concepts, each with 12 images (viewed once), and the testing set selects one unseen image from 200 of these concepts, repeated 12 times. The THINGS-fMRI training set comprises 3 subjects, 720 concepts, each associated with 12 images (viewed once), and the testing set includes 100 concepts, with a single image selected from each and repeated 12 times. To facilitate zero-shot evaluation, we reclassified the data during experiments (See Appendix for more details).
Table 1. Multimodal Retrieval Performance. We present the retrieval performance of various methods on the THINGS-EEG2, THINGS-MEG and THINGS-fMRI datasets. Retrieval refers to the task of correctly identifying paired neural latent representations from the 200 or 100 class test set clip embeddings. Performance is evaluated across joint subject retrieval tasks, and we report results for several retrieval configurations: 2-way, 4-way, 10-way, as well as Top-1 and Top-5 accuracy for the 200-way retrieval task (100 for fMRI). BrainFLORA is compared with previous SOTA methods, including NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ), UMBRAE (xia2024umbrae, ), ATM-S (li2024visual, ), CognitionCapturer (zhang2024cognitioncapturer, ), MindBridge (wang2024mindbridge, ), WAVE (wang2024decoding, ) and MB2C (wei2024mb2c, ). In-subject results refers to Appendix.
Model THINGS-EEG2 THINGS-MEG THINGS-fMRI 2-way 4-way 10-way 200-way 200-way 2-way 4-way 10-way 200-way 200-way 2-way 4-way 10-way 100-way 100-way Top-1 Top-1 Top-1 Top-1 Top-5 Top-1 Top-1 Top-1 Top-1 Top-5 Top-1 Top-1 Top-1 Top-1 Top-5 NICE 90.10 77.15 57.15 13.80 35.80 60.62 39.62 20.87 2.50 7.88 87.00 74.67 55.00 22.00 51.00 EEGNetV4 88.15 77.05 59.10 12.15 35.85 71.13 48.63 29.38 3.88 11.88 86.00 74.67 55.33 18.00 44.67 B.D. 90.05 76.85 55.85 12.45 34.15 62.12 35.50 17.63 1.88 6.50 72.67 58.33 38.00 6.67 26.00 MindEyeV2 90.25 79.85 64.30 17.35 43.65 64.25 40.25 20.37 2.13 8.38 92.00 80.33 64.33 21.33 60.33 UMBRAE 66.50 41.95 21.85 2.90 10.00 61.75 35.50 16.63 2.00 6.50 85.67 70.67 50.67 16.67 42.67 ATM-S 93.35 83.70 68.60 19.70 48.60 77.00 59.25 38.88 6.25 20.12 92.67 81.67 65.33 24.00 55.33 CognitionCapturer 92.90 84.00 67.85 20.50 50.70 67.88 47.63 23.57 2.38 10.37 86.67 73.00 55.33 20.67 48.00 MindBridge 89.10 75.35 57.45 15.00 39.15 64.88 40.12 20.13 1.87 7.37 88.33 73.33 55.67 16.67 54.33 WAVE 89.40 75.85 58.80 15.00 39.05 67.50 44.75 24.00 3.00 11.38 88.67 71.67 52.33 20.33 45.33 MB2C 85.20 70.25 49.65 10.20 29.85 61.75 33.25 17.00 1.25 5.50 84.67 69.00 49.67 18.67 42.33 BrainFLORA-uni 95.55 86.90 73.45 25.35 57.30 81.75 64.50 46.62 8.00 24.38 91.33 79.00 63.00 26.33 57.33 BrainFLORA-multi 94.05 87.30 73.15 25.05 56.35 80.50 61.88 39.75 6.88 23.38 92.33 84.67 70.67 28.33 63.33
Table 2. Multimodal Visual Reconstruction. We quantitatively evaluated reconstruction quality across EEG, MEG, and fMRI modalities, reporting averaged results from all subjects. Unlike previous works that perform training and inference within individual subjects, BrainFLORA operates in a cross-subject, cross-modal setting for both training and reconstruction evaluation. Additional evaluation details can be found in Appendix Sec.B.
Dataset Low-level High-level
[…]
THINGS-EEG2 0.101 0.360 0.653 0.708 0.607 0.635 0.819 0.625
THINGS-MEG 0.074 0.261 0.545 0.595 0.553 0.586 0.818 0.688
THINGS-fMRI 0.079 0.348 0.595 0.622 0.604 0.576 0.812 0.661
4.1.2. Implementation and Training Setup.
All experiments, including both training and inference procedures across modalities, were conducted using two NVIDIA H20-96GB GPU. For visual encoding, we employed CLIP (ViT-L-14) to generate image embeddings. Each modality encoder was trained on the original training set of the THINGS dataset for 50 epochs, with a learning rate of 3e-4 and a batch size of 360, utilizing the AdamW optimizer. During the forward pass, batches from different modalities were processed sequentially, with each batch maintaining a size equal to the specified batch size. Hyper-parameters were shared across all modalities.
4.2. Results on Retrieval
Multimodal retrieval
Image retrieval metrics are used to quantify the extent of visual information captured in neural embeddings. The effectiveness of image retrieval task is evaluated by computing the cosine similarity between embeddings derived from EEG, MEG, and fMRI signals. We trained the model for 150 epochs under a joint-subject setting across EEG, MEG, and fMRI data, optimizing only the contrastive loss (Sec. 3.2) to evaluate representation alignment of semantic representations during retrieval. And Tab. 1 presents the average retrieval performance across all subjects for EEG, MEG and fMRI. Under challenging mixed-modality training/testing conditions, our approach establishes new SOTA results on the THINGS joint-subject retrieval benchmark, demonstrating consistent improvements across all three neural modalities (EEG: Top-1 25.05%, MEG: Top-1 6.88%, fMRI: Top-1 28.33%). We observe degraded THINGS-MEG performance ($\mathrm{\Delta}$Acc = -48.78% vs original splits, see Appendix for more details) due to our controlled evaluation protocol: test images now contain single-trial neural responses only, eliminating the signal averaging advantage present in standard benchmarks.
Table 3. Brain Captioning. We train and evaluate brain captioning performance across three modalities independently (THINGS-EEG2, THINGS-MEG, and THINGS-fMRI) under a joint cross-subject training and evaluation setting.
Dataset BLEU1 $\uparrow$ BLEU2 $\uparrow$ BLEU3 $\uparrow$ BLEU4 $\uparrow$ METEOR $\uparrow$ ROUGE $\uparrow$ CIDEr $\uparrow$ SPICE $\uparrow$ CLIP-S $\uparrow$ RefCLIP-S $\uparrow$
THINGS-EEG2 36.36 30.80 25.76 19.67 20.01 45.86 2.86 15.62 42.74 51.80
THINGS-MEG 36.53 31.01 25.99 19.88 19.79 46.01 3.15 15.51 43.01 51.55
THINGS-fMRI 37.96 31.59 26.29 20.11 20.03 46.69 4.40 15.55 44.21 52.96
* The evaluation metrics for BrainFLORA were computed using the BrainHub (xia2024umbrae, ) benchmark.
4.3. Results on Reconstruction
Multimodal reconstruction
While recent work has achieved breakthroughs in fMRI-to-image reconstruction using diffusion models (takagi2023high, ; chen2023seeing, ; scottimindeye2, ), we present the first unified framework capable of reconstructing images from multimodal neural data (EEG/MEG/fMRI) via a single model. Our approach extends MindEye2 (scottimindeye2, ) by integrating SDXL (podell2023sdxl, ) with IP-Adapter (ye2023ip, ), while introducing two key innovations: (1) a shared cross-modal encoder that preserves modality-specific features through contrastive alignment, and (2) a lightweight prior diffusion (ramesh2022hierarchical, ) that maps neural embeddings to the SDXL latent space. Similar to (li2024visual, ), our method enables multi-level controllable generation through hierarchical conditioning. As shown in Tab. 2, this pipeline achieves impressive reconstruction quality while eliminating the need for modality-specific decoders. Quantitative analysis reveals statistically insignificant differences between BrainFLORA-uni and BrainFLORA-multi ($\mathrm{\Delta}$Acc=7.58%, Tab. 1), demonstrating that the added complexity of modality-specific prior diffusion training yields no measurable advantage for cross-modal retrieval.
4.4. Results on Captioning
Multimodal caption
Building on evidence that latent-space reconstruction outperforms image-to-text pipelines (li2024realmind, ), we assessed our model’s captioning capability by decoding unified neural embeddings via OpenFlamingo (awadalla2023openflamingo, ). We identify a critical stability trade-off in multimodal reconstruction: while concurrent CLIP and MSE loss optimization improves feature alignment in prior work (benchetrit2024brain, ; li2024visual, ), our experiments demonstrate catastrophic interference effects ($\mathrm{\Delta}$ loss ¿ 100%) during training. This motivates our MSE-only training paradigm. During the inference phase of OpenFlamingo, we selected an in-distribution image (category: cat) from the training set and used the prompt ”An image of” as one-shot text to guide the neural signal for caption generation. As shown in Tab. 3, single-modality evaluations (it was adopted due to GPU memory constraints) achieve competitive performance with SOTA neural decoding methods. Multimodal fusion results remain an open challenge for future work.
Figure 4. Similarity analysis of the concept representation between EEG, MEG and fMRI across all subjects. a-c: We calculate the representation similarity matrix (RSM) between samples. RSMs for the test set consisting of all objects, created based in the projected BrainFLORA embedding (left). Correlation between the predicted and measured similarity on all object pairs (middle). The distribution of predicted and measured concept embedding are shown (right).
4.5. In-depth Analysis
To identify the distribution of concept representations related to visual perception across different neural modalities in the brain, we used the projection layer in CoCoG (wei2024cocog, ) to map neural modality embeddings into a 42-dimensional concept space, obtaining concept embeddings for each modality.
4.5.1. Concept Representation Similarity Analysis.
To assess the fidelity of the resulting embedding, we conducted an evaluation by exhaustively gathering concept embedding for a group of all objects not included in the BrainFLORA’s training data. By computing their similarity, we compared this measured similarity matrix with the one predicted by the model’s embedding in Fig. 4. Significant positive correlations were observed between the model-predicted and measured Representational Similarity Matrices (RSMs) (correlation coefficients: 0.50 for EEG, 0.56 for MEG, and 0.40 for fMRI). These results validate that the 42-dimensional embeddings we analyzed closely reflect the similarity space underlying the corresponding visual concept embedding.
4.5.2. Concept Embedding Retrieval.
We show the forward and the backward retrieval performance of image concept embedding using predicted embedding from three different neural modalities in Fig. 5. These results indicate that successful cross-modal retrieval performance depends more on the discriminative quality of cross-modal alignment rather than on on the fidelity of preserved intra-modal similarity structures. This finding is consistent with principles established in contrastive learning research, where optimizing for cross-modal mutual information has been shown to be more crucial than maintaining intra-modal structural relationships.
4.5.3. Emergent Object Category Information.
We visualize the global structure of the acquired embeddings through a multidimensional scaling (MDS)-initialized t-SNE plot representing all objects in Subject-01 across different neural modalities in Fig. 6. Objects with similar embedding dimensional values appear proximally in the visualization, demonstrating that items belonging to the identical semantic category cluster together consistently across EEG, MEG, and fMRI modalities. We provide t-SNE visualizations of each modality’s embedding alongside the image modality in the appendix (Fig.3), demonstrating their consistent distributions after dimensionality reduction.
5. Discussion and Conclusion
We present BrainFLORA, the first unified model that establishes universal alignment of visual representations across three major neuroimaging modalities—EEG, MEG, and fMRI—via a shared latent space. At its core, BrainFLORA introduces a universal projector that bridges domain-specific neural signatures while preserving modality-invariant semantic features.
Figure 5. The forward and backward retrieval performance between EEG, MEG and fMRI.
Our framework enables two previously underexplored capabilities: (1) cross-modal representation translation without paired neural data, and (2) multimodal concept analysis within a single end-to-end architecture. Rigorous experiments on THINGS-based datasets demonstrate SOTA performance in neural decoding accuracy (zero-shot in EEG, MEG) compared to modality-specific baselines. Beyond empirical results, we provide theoretical insights into the geometric properties of the learned shared space. Through extensive empirical studies, we demonstrate that EEG, MEG, and fMRI signals—when aligned via cross-modal training—naturally converge to concept representation structures that mirror the hierarchical organization of physical-world objects. Crucially, this emergent geometric consistency is automatically learned through modality alignment without explicit supervision.
Figure 6. The t-SNE visualization of test set objects in Subject-01 from each neural modality, showing emergent clusters in the learned concept embedding between EEG, MEG and fMRI.
Our framework has several limitations that warrant further investigation. First, the uneven distribution of samples across modalities and classes hinders the framework’s ability to learn modality-specific representations. Future work could explore balanced pre-training strategies or reweighting schemes. Second, Shared parameters cause suboptimal performance in tasks like reconstruction and caption generation compared to specialized SOTA methods. This gap could potentially be addressed by leveraging larger datasets and more sophisticated model architectures. Finally, The THINGS dataset’s limited size and category overlap between train/test splits (e.g., fMRI’s inclusion relationships) introduce bias. Larger-scale cross-modal datasets with rigorous category splits are needed for robust evaluation.
Despite these limitations, BrainFLORA establishes the feasibility of multimodal grand unified neural decoding through effective cross-modal complementary information integration, achieving consistent performance across diverse tasks. Future work may focus on two key directions: (i) using larger neural datasets, both paired and unpaired, to train robust multimodal foundation models, and (ii) leveraging cutting-edge model architectures that incorporate advanced components for representation, alignment, and fusion, tailored to diverse neural decoding tasks. With such a unified framework, more scientific questions can be studied, such as how these semantic representations across modalities are unified and adapted to personal experiences and knowledge.
These advances not only push the boundaries of neural decoding but also shed light on the intricate interactions between diverse neural data modalities. Moving forward, two critical directions emerge: (1) leveraging larger-scale neural datasets—both paired and unpaired—to train robust multimodal foundation models, and (2) pioneering grand unified architectures that integrate representation learning, modality alignment, and fusion mechanisms for generalizable neural decoding.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62472206), Shenzhen Science and Technology Innovation Committee (RCYX20231211090405003, KJZD20230923115221044), Guangdong Provincial Key Laboratory of Advanced Biomaterials (2022B1212010003), and the open research fund of the Guangdong Provincial Key Laboratory of Mathematical and Neural Dynamical Systems, the Center for Computational Science and Engineering at Southern University of Science and Technology.
Supplementary Material:
BrainFLORA: Uncovering Brain Concept Representation via Multimodal Neural Embeddings
Appendix A Appendix Overview
  1. Sec. B: Additional Ablation Experiments.
  2. Sec. C: Additional Implementation Details.
  3. Sec. D: Evaluation Details.
  4. Sec. E: Comparison with Prior Works.
  5. Sec. F: Additional Qualitative Results.
Appendix B Additional Ablation Experiments
B.1. Ablation Study of BrainFLORA
To assess the contribution of each component within our framework, we conducted ablation studies on various model modules in Tab. 4. Using Medformer (wang2024medformer, ) as the baseline, we evaluated the impact of the temporal-spatial convolution module and the Mixture of Experts (MoE) module. Additionally, we explored replacing the MoE module with the Perceiver as the universal projection module; however, this approach resulted in a prohibitively large number of parameters, slower computation times, and suboptimal performance.
The experimental results substantiate that our proposed architecture effectively optimizes the trade-off between model capacity and computational efficiency. The temporal-spatial convolutional module preserves predictive performance while substantially reducing the parameter count, thereby facilitating the scalable processing of high-dimensional neural datasets. More critically, our soft-routing universal projection mechanism, when trained with heterogeneous multimodal neural data, learns invariant and transferable feature representations. Traditional MoE module selects a single expert based on the maximum score, while softrouter implements a weighted combination of experts. This architectural paradigm is particularly advantageous for joint-subject modeling, as the exposure to diverse neurophysiological patterns across modalities during training enhances BrainFLORA’s robustness in addressing inter-subject variability while maintaining a significantly lower computational burden compared to alternatives such as Perceiver (alayrac2022flamingo, ). The empirical results demonstrate that this routing strategy in universal projection facilitates knowledge transfer across modalities, contributing to strong generalization capabilities in zero-shot cross-modal inference.
The ablation study corroborates our design rationale, highlighting the efficacy of multimodal training in bolstering model generalization—an attribute that becomes increasingly pivotal when scaling to expansive neural datasets encompassing extensive subject pools and multiple trial repetitions.
Figure 7. Effect of data size in training and test set (THINGS-EEG2), and weakly-supervised subject adaptation (THINGS-EEG1).
B.2. Supervised Scaling Up on Data Size
We analyzed the effect of training set size on joint subject training performance using THINGS-EEG2, as illustrated in 7A. This factor is critical for model efficacy. The training set comprised $1654$ distinct concepts $\times$ $10$ image conditions, each repeated four times. We preserved all four repetitions for training. Expanding the dataset size by increasing the number of conditions yielded substantial improvements in decoding accuracy, particularly in transitions from $20\%$ to $40\%$ and from $40\%$ to $60\%$. Additionally, augmenting repetitions contributed significantly to performance gains, notably from $80\%$ to $100\%$. These findings indicate that, under joint-subject training, larger datasets facilitate further performance enhancements. Our approach demonstrates strong scalability, underscoring its potential for improving neural decoding with increased data availability.
[…]
Figure 8. Relative improvement (%) in EEG/MEG/fMRI cross-modal retrieval accuracy under joint-subject training compared to in-subject baselines, evaluated across methodological approaches.
Table 4. Retrieval performance of BrainFLORA under varying input projection mechanisms. The first two rows compare time-series Transformers trained using an attention-based algorithm without any projection or feature fusion. The following rows illustrate the impact of BrainFLORA enhancements. “Med.”: A multi-granularity time-series model Medformer (wang2024medformer, ) designed for medical data. “Med. (TSConv)”: Medformer with the addition of temporal-spatial convolution. “Re. EEG”: Retrieval performance on the THINGS-EEG2 dataset. “Re. MEG”: Retrieval performance on the THINGS-MEG dataset. “Re. fMRI”: Retrieval performance on the THINGS-fMRI dataset.
# Description Para. Model Uni. Proj. Softrouter Re. EEG $\uparrow$ Re. MEG $\uparrow$ Re. fMRI $\uparrow$
1 Med. (Linear) 97.9M Time series 18.15 3.50 25.00
2 Med. (TSConv) 18.67M BrainFLORA-uni 26.50 3.00 24.00
3 +Uni. Proj. 162.0M BrainFLORA-multi 25.05 6.88 28.33
4 +Softrouter 162.0M BrainFLORA-multi 26.45 8.38 24.33
Appendix C Additional Implementation Details
C.1. Neural Modality Encoders
EEG Encoder. The shape of an EEG signal is $R^{63\times 250}$. With Medformer, we get multiple sets of tokens output $x\in\mathbb{R}^{221\times 250}$. By temporal-spatial convolution layers, we then resize the tensor $x\in\mathbb{R}^{221\times 250}$ into a 1D tensor $x\in\mathbb{R}^{1\times 1024}$.
MEG Encoder. The shape of an MEG signal is $R^{271\times 201}$. With Medformer, we get multiple sets of tokens output $x\in\mathbb{R}^{178\times 250}$. By temporal-spatial convolution layers, we then resize the tensor $x\in\mathbb{R}^{178\times 250}$ into a 1D tensor $x\in\mathbb{R}^{1\times 1024}$.
fMRI Encoder. In the THINGS-fMRI data set, the original voxel number of visual ROI (Region of Interest) of each of the three subjects was 6036, 5944 and 5238. In order to facilitate model processing, pad was unified to 7000 when data was loaded. So the shape of an fMRI signal is $R^{7000}$. We tokenize it with a linear layer: $\text{Linear}(C_{\text{in}}=7000,C_{\text{out}}=8192)$. We then resize the output tensor $x\in\mathbb{R}^{8192}$ into a 2D tensor $x\in\mathbb{R}^{8\times 1024}$ to align with the input of the transformer encoder. With Medformer, we get multiple sets of tokens output $x\in\mathbb{R}^{899\times 250}$. By temporal-spatial convolution layers, we then resize the tensor $x\in\mathbb{R}^{899\times 250}$ into a 1D tensor $x\in\mathbb{R}^{1\times 1024}$.
Figure 9. t-SNE visualization of concept embedding predicted by BrainFLORA in Subject- 01 between EEG, MEG and fMRI.
C.2. Low-Level Pipeline
Compared to vision-centric pretraining paradigms such as ViT, ResNet, and DINO, the CLIP vision model exhibits a deficiency in capturing fine-grained low-level visual features. To mitigate this limitation, our framework integrates a dedicated low-level visual reconstruction pipeline. Specifically, we aim to reconstruct fundamental perceptual attributes—such as contour, posture, and orientation—by leveraging EEG-derived representations. This is achieved through an alignment mechanism with the latent space of a variational autoencoder (VAE), facilitating the recovery of pixel-level structural information. By enforcing this latent alignment, our approach enhances the model’s capacity to infer and preserve crucial low-level visual semantics from neural signals, thereby improving the fidelity of EEG-to-vision mappings.
We trained the low-level reconstruction pipeline for 200 epochs, employing a latent mean squared error (MSE) loss in conjunction with a contrastive learning loss and a variational autoencoder (VAE) image reconstruction loss. The objective was to align the $4\times 64\times 64$ EEG-derived latent representations—obtained via a projection layer and an upsampling CNN—with the VAE latent space. However, we observed that the reconstruction loss and contrastive learning loss underperformed compared to solely optimizing the latent space loss, while also imposing significantly higher GPU memory demands.
Furthermore, our investigation revealed that incorporating a low-level visual model for knowledge distillation in the low-level reconstruction pipeline not only failed to enhance VAE latent alignment but also exacerbated overfitting. These findings suggest that zero-shot low-level visual reconstruction from EEG signals lacks stability and may introduce misleading artifacts in the generated outputs.
In our framework, when utilizing the low-level reconstruction pipeline, we typically set the inference steps of SDXL to 10 (or SDXL-Turbo to 4) and configure the image-to-image denoising strength to 0.5.
Appendix D Evaluation Details
In this section, we will give more evaluation details.
Visual Retrieval tasks.
We use forward retrieval (all image embeddings in test set is retrieved using one neural data embedding) for evaluation. For the neural data of each modality, we output it as $x\in\mathbb{R}^{1\times 1024}$ and calculate the dot product similarity with the image embedding, taking the image with the greatest similarity as the image output by zero-shot. For THINGS-EEG1, the test set has 12 conditions per concept, so we only take the first image condition among the 200 concepts in the test set, that is, we use 200 images from 200 categories to test. For THINGS-EEG2, the test set has only 1 condition per concept, so we test with 200 images from 200 categories. For THINGS-MEG, there are 200 concepts in the test set, each of which has 12 conditions, so we only use the first condition of each concept as the test, with a total of 200 images from 200 categories to test. For THINGS-fMRI, the test set is 100 concepts, where each concept has 1 condition, so a total of 200 images from 200 categories are tested.
Visual Reconstruction tasks.
The test set configuration of the visual reconstruction task is exactly the same as that of the retrieval task, except that the image reconstruction has more indicators to measure the performance of the image reconstruction. It should be emphasized that, in contrast to existing approaches on the THINGS dataset which train models independently for each subject’s neural data and conduct visual reconstruction, BrainFLORA-multi necessitates addressing highly heterogeneous multimodal and cross-subject neural data in both the training phase and visual reconstruction inference evaluation. Moreover, within the fMRI modality, BrainFLORA employs the THINGS-fMRI dataset, which is smaller in scale than the NSD dataset used in existing major works, thus requiring the model to more efficiently utilize limited data resources.
Visual Captioning tasks.
The idea of employing Prior diffusion to alter the distribution of model features is inspired by the two-stage image reconstruction approach described in ATM-S (li2024visual, ), with the expectation of achieving similar effects in the domain of caption generation. Although the final quantitative results show only marginal differences, the incorporation of Prior diffusion leads to more structurally coherent and semantically accurate captions that better reflect the content of the original images. The test set configuration for the visual captioning task is exactly the same as for the retrieval task. The THINGS image data set lacks the original text captions annotation. To generate captions for the test set images as ground truth, we used the prompt ”Please provide a concise and formal sentence to describe the image, beginning with ’An image of,’ and keeping it to approximately 10 words.” from the large model Kimi to obtain descriptions for each image in the test set. By using different MMLMs as interfaces, we can generate captions of different qualities from the latent, and we give an assessment of captioning quality for different MMLMs adapted to BrainFLORA.
Appendix E Comparison with Prior Works
The key distinction between BrainFLORA and previous neural decoding approaches lies in its demonstration that specialized neural encoders, when coupled with a universal projection layer, can effectively align diverse neural modalities with visual representations in a unified framework. While prior works typically develop independent architectures for each modality, BrainFLORA achieves multimodal integration across EEG, MEG, and fMRI signals while maintaining a compact parameter footprint (7.36M). The empirical results demonstrate competitive performance across modalities compared to modality-specific models, while ablation studies (Tab. 4) reveal that joint training within this unified framework provides particular benefits for modalities with limited data availability.
The effectiveness of BrainFLORA’s architecture in handling diverse neural signals suggests broader implications for scalable neural decoding. By successfully integrating multiple modalities through a shared projection space while maintaining computational efficiency, BrainFLORA provides insights into architectural design principles for larger-scale neural models. The framework’s demonstrated capabilities in addressing fundamental challenges - from cross-subject variability to zero-shot generalization - while preserving modality-specific characteristics, establish a foundation for developing more comprehensive neural decoding systems. These results not only validate the feasibility of unified multimodal frameworks but also suggest promising directions for scaling neural decoding models toward more ambitious multimodal applications.
Appendix F Additional Qualitative Results
F.1. Additional Retrieval Results
To achieve cross-modal concept-level zero-shot retrieval, we realigned the THINGS-MEG dataset split to match THINGS-EEG2. Tab.5 demonstrates the impact of this realignment and trial averaging on retrieval performance. Notably, due to the realignment, some concepts in both training and test sets inherit different numbers of trials from the original MEG dataset structure. To maintain consistency, we standardized to one trial per concept in the unave. (unaveraged) setting. However, for concepts that were originally in the MEG test set (which had 12 trials per concept), we also evaluated a ”averaged” setting where these 12 trials were averaged. The performance improvement from BrainFLORA-unimodal (unave.) to BrainFLORA-unimodal demonstrates the benefit of this trial averaging, which effectively improves the signal-to-noise ratio for these concepts.
Table 5. Impact of average retrieval performance across multiple conditions on the THINGS-MEG test set. We report the retrieval performance of BrainFLORA on the THINGS-MEG dataset under various evaluation settings.
Model
2-way
Top-1
4-way
Top-1
10-way
Top-1
200-way
Top-1
200-way
Top-5
BrainFLORA-unimodal (unave.) 72.38 54.62 36.88 5.75 17.87
BrainFLORA-unimodal 81.75 64.50 46.62 8.00 24.38
BrainFLORA-multimodal (unave.) 74.75 52.00 36.38 6.12 18.25
BrainFLORA-multimodal 80.50 61.88 39.75 6.88 23.38
The original THINGS-MEG dataset presents this evaluation scenario in its data partitioning: the training set contains 1,654 concepts (12 images per concept, single repetition), while the test set comprises 200 concepts (1 image per concept, 12 repetitions per image) in a concept-level zero-shot setting. In this experiment, BrainFLORA-unimodal was trained without the MoE-based universal projection and multimodal training, using an independent model for each subject’s MEG recordings in an in-subject evaluation protocol. While our model has demonstrated superior performance in cross-subject zero-shot scenarios with richer data availability, it maintains competitive performance in this more constrained setting, as evidenced in Tab.6. The results validate that BrainFLORA’s core architecture provides robust performance even when trained on single-subject data with limited per-concept samples, without leveraging the benefits of multimodal training data and cross-subject information.
Table 6. Retrieval performance on the original THINGS-MEG dataset. We trained and evaluated various methods on the original THINGS-MEG training and test sets. Each method was independently trained and assessed for each subject, with results reported for different retrieval configurations: 2-way, 4-way, 10-way, as well as Top-1 and Top-5 accuracy for the 200-way retrieval task.The methods compared include NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D., MindEyeV2 (scottimindeye2, ) and ATM-S (li2024visual, ).
Model
2-way
Top-1
4-way
Top-1
10-way
Top-1
200-way
Top-1
200-way
Top-5
NICE 89.29 76.94 60.19 17.12 40.15
EEGNetV4 85.34 71.78 52.85 12.76 32.38
B.D. 77.16 57.76 37.82 6.06 19.38
ATM-S 90.25 79.50 63.66 17.84 44.73
BrainFLORA-unimodal 87.75 75.25 57.50 15.62 39.37
Table 7. In-subject Retrieval on THINGS-EEG2. We report the retrieval performance of different methods on the THINGS-EEG2 (in-subject) datasets. Each method is trained and evaluated in each subject independently, with results for different retrieval configurations: 2-way, 4-way, 10-way, and the Top-1 and Top-5 accuracy of 200-way. The methods compared include NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ), UMBRAE (xia2024umbrae, ), ATM-S (li2024visual, ), CognitionCapturer (zhang2024cognitioncapturer, ), MindBridge (wang2024mindbridge, ), WAVE (wang2024decoding, ) and MB2C (wei2024mb2c, ).
Model THINGS-EEG2 (in-subject)
2-way 4-way 10-way 200-way 200-way
Top-1 Top-1 Top-1 Top-1 Top-5
NICE 93.23 83.93 69.22 21.67 51.34
EEGNetV4 91.42 80.21 63.37 16.84 42.58
B.D. 89.03 56.77 56.77 13.29 35.50
CogCap 93.15 82.85 69.35 22.05 51.60
MB2C 78.40 62.25 43.75 8.85 25.20
Wave 88.65 73.70 54.70 13.00 34.50
MindBridge 89.10 75.35 57.45 15.00 39.15
ATM-S 94.70 86.73 74.00 26.85 57.21
MindEyeV2 92.50 82.80 66.10 23.80 50.25
UMBRAE 69.75 49.15 28.80 3.50 12.25
BrainFLORA-unimodal 93.75 85.70 69.65 21.85 51.65
Our evaluation demonstrates that BrainFLORA achieves strong performance in both in-subject and cross-subject settings on the THINGS-EEG2 dataset. When tested in the in-subject paradigm, where models are independently trained and tested on individual subjects’ data as is shown in Tab.7, BrainFLORA maintains competitive performance.
Notably, our cross-subject evaluation results show better retrieval performance compared to the in-subject setting, validating our model’s effectiveness in addressing cross-subject challenges. BrainFLORA’s robust performance can be attributed to its efficient encoder architecture across different data scales and the flexibility to incorporate or remove components like the MoE module. These results confirm BrainFLORA’s capability to deliver strong performance across various experimental paradigms while effectively leveraging larger-scale datasets when available.
Table 8. In-subject Retrieval on THINGS-MEG. We report the retrieval performance of different methods on the THINGS-MEG dataset in the in-subject setting. Each method is trained and evaluated in each subject independently, with results for different retrieval configurations: 2-way, 4-way, 10-way, and the Top-1 and Top-5 accuracy of 200-way. The methods compared include NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ) and ATM-S (li2024visual, ).
Model THINGS-MEG (in-subject)
2-way 4-way 10-way 200-way 200-way
Top-1 Top-1 Top-1 Top-1 Top-5
NICE (unave.) 68.65 43.63 24.88 2.63 9.25
EEGNetV4 (unave.) 69.25 49.63 29.00 4.37 13.25
B.D. (unave.) 56.75 35.00 16.12 0.75 5.37
MindEyeV2 (unave.) 68.12 45.13 24.62 2.25 9.37
ATM-S (unave.) 80.13 60.75 42.75 7.38 21.75
BrainFLORA-unimodal(unave.) 76.25 59.12 37.25 6.25 17.25
NICE 79.75 60.75 40.00 6.00 20.50
EEGNetV4 80.37 62.87 41.00 6.75 21.50
B.D. 64.25 43.32 21.50 2.00 8.50
MindEyeV2 76.25 56.88 37.62 6.00 18.75
ATM-S 82.87 68.00 46.13 8.12 27.12
BrainFLORA-unimodal 81.75 66.00 48.13 10.25 26.25
The in-subject evaluation on the THINGS-MEG dataset demonstrates that BrainFLORA-unimodal achieves state-of-the-art performance in the averaged condition while maintaining robust performance in the unaveraged setting (Tab.8). This superior performance, particularly with high-quality averaged MEG signals, validates BrainFLORA’s strong decoding capabilities even when operating on smaller-scale datasets.
This consistent performance can be attributed to two fundamental aspects of our architecture: first, the model’s adaptable design that effectively accommodates various experimental paradigms, and second, its ability to efficiently utilize the enhanced signal-to-noise ratios in averaged MEG data without relying on the advantages of large-scale datasets. These findings further substantiate BrainFLORA’s effectiveness as a versatile neural decoding framework, demonstrating robust performance across both joint-subject and in-subject experimental paradigms.
Table 9. In-subject Retrieval on THINGS-fMRI. We report the retrieval performance of different methods on the THINGS-fMRI dataset. Each method is trained and evaluated in each subject independently, with results for different retrieval configurations: 2-way, 4-way, 10-way, and the Top-1 and Top-5 accuracy of 200-way. The methods compared include NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ), UMBRAE (xia2024umbrae, ), ATM-S (li2024visual, ), CognitionCapturer (zhang2024cognitioncapturer, ), MindBridge (wang2024mindbridge, ), WAVE (wang2024decoding, ) and MB2C (wei2024mb2c, ).
Model THINGS-fMRI (in-subject)
2-way 4-way 10-way 100-way 100-way
Top-1 Top-1 Top-1 Top-1 Top-5
NICE 90.00 82.67 62.67 25.00 56.00
EEGNetV4 89.00 76.00 56.00 20.33 50.67
B.D. 80.33 69.67 33.33 9.67 24.00
CogCap 93.67 81.33 64.00 27.33 64.00
MB2C 93.00 80.33 61.33 23.67 57.00
Neuro.V2L 86.67 70.67 50.67 15.00 39.67
Wave 91.00 79.67 65.67 29.00 61.00
ATM-S 95.00 81.00 66.67 25.67 60.33
MindEyeV2 95.00 86.67 76.67 32.67 69.00
BrainFLORA-unimodal 92.33 87.00 67.00 24.33 61.67
The in-subject evaluation on the THINGS-fMRI dataset demonstrates BrainFLORA-unimodal’s strong adaptability in decoding capabilities, even when operating on limited single-subject data. As shown in Tab.9, our model maintains competitive performance, validating its effectiveness across varying data scales and modalities. These results substantiate that BrainFLORA, while primarily optimized for cross-subject scenarios with larger-scale datasets, retains robust decoding capabilities even when constrained to limited single-subject fMRI data. This finding further underscores the versatility of our architectural design across diverse experimental paradigms.
Examining in-subject retrieval performance, Tab. 7, Tab. 8 and Tab. 9 demonstrate BrainFLORA’s consistent performance across EEG, MEG, and fMRI modalities in in-subject settings, showing that our model, while optimized for cross-subject generalization, maintains competitive performance when trained on single-subject data. Notably, while traditional approaches often suffer significant performance degradation in cross-subject scenarios, BrainFLORA maintains or improves its performance when moving from in-subject to joint-subject settings, proposing a solution to joint-subject large neural models.
F.2. Additional Concept Analysis
We adopted a methodology similar to that presented in (shen2024neuro, ) for localizing semantic concepts within neural signals. Specifically, we employed Honeybee (cha2024honeybee, ) to extract the target concepts from natural language with the prompt ”Describe the main concept in the picture in three words”. These concepts were encoded by the CLIP ViT-L-14 text encoder and used as target representations for integrated Grad-CAM, facilitating the spatial localization of these concepts within brain signals. We trained a unified encoder incorporating all three input modalities. The final layers of this model were leveraged to extract semantic features. Fig. 10 visualizes the localization outcomes on the THINGS-fMRI dataset, showing the discrimination of various semantic information within brain signals in response to novel visual stimuli.
We conducted ablation studies on the localized semantic concepts. After identifying the relevant concepts within the original fMRI voxels from visual cortex, we selectively masked the corresponding voxels. Feature extraction and visual reconstruction were subsequently performed using these modified brain signals. In Fig. 11, the targeted removal of fMRI voxels from specific brain regions associated with particular semantic concepts led to the exclusion of these semantics in the reconstructed visual representation.
Figure 10. Gradient heatmaps of brain activity generated by BrainFLORA. It shows the neural representation of different semantic information in the brain in response to the same visual stimulus, such as the semantics ’green’ and ’tall’ from the bamboo image.
Figure 11. Concept analysis on THINGS-fMRI using BrainFLORA through semantic signal nullification. We show the impacts on visual reconstruction by ablating certain concepts from visual stimuli.
F.3. Additional Reconstruction Results
We present a qualitative comparison of the highest-, moderate-, and lowest-fidelity generated images in Fig. 12. EEG data were randomly sampled from all subjects viewing 200 images, and corresponding EEG-derived embeddings were extracted to condition the image generation process. By computing the cosine similarity between the CLIP embeddings of the generated and original images, we identified six examples for each fidelity category. In the highest-fidelity group, the generated images exhibit strong semantic correspondence with the original images while effectively preserving low-level visual attributes. In the moderate-fidelity group, the generated images maintain semantic coherence with the original stimuli, yet exhibit partial alterations in low-level visual details. In the lowest-fidelity group, both semantic integrity and low-level visual structures are substantially distorted, leading to a pronounced divergence from the original images. See Fig. 13 and Fig. 14 for MEG and fMRI reconstruction results.
Figure 12. Examples of cross-modal EEG-guided cross-subject visual reconstruction from BrainFLORA. From top to bottom, we exhibit the best, median, and worst 6 generated images, respectively. We show the images subjects seen and the generated images by our two-stage image generator.
Figure 13. Examples of cross-modal MEG-guided cross-subject visual reconstruction from BrainFLORA. From top to bottom, we exhibit the best, median, and worst 6 generated images, respectively. We show the images subjects seen and the generated images by our two-stage image generator.
Figure 14. Examples of cross-modal fMRI-guided cross-subject visual reconstruction from BrainFLORA. From top to bottom, we exhibit the best, median, and worst 6 generated images, respectively. We show the images subjects seen and the generated images by our two-stage image generator.